I'm not sure if I have the correct type of question, but I need to know how to solve this problem. It is making me crazy!
5/2x+6 + 46/7x = 2
Please help!

Answer by stanbon(57342)

5/(2x+6) + 46/(7x) = 2
-----
Multiply thru by (7x)(2x+6) to get:
----------------------------
5(7x) + 46(2x+6) = 2(7x)(2x+6)
----
35x + 92x + 46*6 = 28x^2+84x
-----------------
28x^2-43x-276 = 0
---
Solve using the Quadratic Formula
----------------
Cheers,
Stan H.